• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ejensen

    Log entry time 08:00:57 on June 09, 2009

    Entry number 280096

    keyword=Owl shift summary

    Shift crew:

    SL - Eric Jensen TO - Pete Markowitz 3rd - Lamiaa El Fassi

    =======================================

    Events:
    -----

    02:00 - Started new compton and happex runs, performed nmr sweep
    03:40 - Started new compton run
    06:00 - Started new compton and happex runs, performed nmr sweep
    07:00 - MCC took beam away for beam studies


    Production Runs
    --------------


    TARGET: Polarized 3He DIRECTION: Transverse CURRENT: 9 uA ENERGY: 2.4272 GeV


    Left # Events Right # Events Notes

    3383 4M 22255 2.84M
    3384 4M 22256 2.85M
    3385 4M 22257 2.85M
    3386 4M 22258 2.84M
    3387 4M 22259 2.84M Started new happex and compton runs, new nmr sweep
    3388 4M 22260 2.86M
    3389 4M 22261 2.86M Started new compton run
    3390 4M 22262 2.86M
    3391 4M 22263 2.85M
    3392 4M 22264 2.86M
    3393 4M 22265 2.85M
    3394 4M 22266 2.85M Started new happex and compton runs, new nmr sweep
    3395 3.5M 22267 2.5M MCC took beam away for beam studies


    HAPPEX runs
    ----------

    02:00 31501
    06:00 31502


    Compton runs
    -----------

    02:00 17267
    04:00 17268
    06:00 17269


    NMR sweeps
    ----------

    02:00 - 58.50%
    06:00 - 58.05%